# Detection and Characterization of Coordinated Online Behavior: A Survey

### v1

@@ 6. Open challenges and future research directions
Research on coordinated online behavior also faces some ethical dilemmas. The process of collecting and analyzing online data involves monitoring user interactions and behaviors, which can raise concerns about surveillance and consent. The inherent subjectivity of coordinated behavior leads to additional challenges, as certain coordinated actions may be punished on some platforms while being tolerated on others [31], resulting in imbalanced interventions and potentially unfair decisions. Characterizing coordinated actors further complicates matters, as this requires assessing the authenticity and harmfulness of actors who might have strong motivations—such as social and political reasons [3]—for remaining anonymous, raising privacy concerns. Additionally, developing methods to detect coordinated actors presents risks as these tools could be misused to silence certain groups or minorities, thereby selectively threatening freedom of expression. Ultimately, these issues demand that great attention to ethical and normative considerations should be devoted in future research.
@@ 7. Conclusions
Acknowledgements.
This work was supported in part by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; in part by PNRR-M4C2 “FAIR-Future Artificial Intelligence Research”-Spoke 1 “Human-Centered AI,” funded under the NextGeneration EU Program under Grant PE00000013; “SoBigData.it—Strengthening the Italian RI for Social Mining and Big Data Analytics”—Prot. IR0000013. H2020 Research Infrastructures (Grant No. 654024).
Author Contributions
Lorenzo Mannocci, Michele Mazza, Anna Monreale, Maurizio Tesconi and Stefano Cresci designed the research; Lorenzo Mannocci and Michele Mazza performed the review; Lorenzo Mannocci, Michele Mazza, Anna Monreale, Maurizio Tesconi and Stefano Cresci wrote the paper; Anna Monreale, Maurizio Tesconi and Stefano Cresci supervised the work; Maurizio Tesconi provided funding.

### v2

@@ 6. Open challenges and future research directions
Ethical dilemmas. Research on coordinated online behavior also faces some ethical dilemmas. The process of collecting and analyzing online data involves monitoring user interactions and behaviors, which can raise concerns about surveillance and consent. The inherent subjectivity of coordinated behavior leads to additional challenges, as certain coordinated actions may be punished on some platforms while being tolerated on others [31], resulting in imbalanced interventions and potentially unfair decisions. Characterizing coordinated actors further complicates matters, as this requires assessing the authenticity and harmfulness of actors who might have strong motivations—such as social and political reasons [126, 61]—for remaining anonymous, raising privacy concerns. Additionally, developing methods to detect coordinated actors presents risks as these tools could be misused to silence certain groups or minorities, thereby selectively threatening freedom of expression. Ultimately, these issues demand that great attention to ethical and normative considerations should be devoted in future research.